find the equation of a line that passes through the point (1,-4) and is parallel to the line 7x-6y=-4
Answer by checkley75(3666)

7x-6y=-4
-6y=-7x-4
y=-(7x/-6)-(4/-6)
y=7x/6+2/3 this line has a slope of 7/6
So a line parallel and passing through (1,-4) with a slope of 7/6 is
-4=7/6*1+b
-4=7/6+b
b=-4-7/6
b=(-24-7)/6
b=-31/6
y=7x/6-31/6